Wilson's Retail Company is planning a cash budget for the next three months. Estimated sales revenue is as follows:

Month Sales Revenue Month Sales Revenue
January $300,000 March $200,000
February 205,000 April 190,000

All sales are on credit; 60 percent is collected during the month of sale, and 40 percent is collected during the next month. Cost of goods sold is 70 percent of sales. Payments for merchandise sold are made in the month following the month of sale. Operating expenses total $44,000 per month and are paid during the month incurred. The cash balance on February 1 is estimated to be $40,000.

Prepare monthly cash budgets for February, March, and April.

Use negative signs only with beginning and ending cash balances, when appropriate. Do not use negative signs with disbursement answers.
Answer #1

Answer-

WILSON RETAIL COMPANY
CASH BUDGET
PARTICULARS FEBRUARY MARCH APRIL
$ $ $
Opening cash balance 40000 29000 43500
Cash collection:-
January sales revenue 300000*40%= 120000
February sales revenue 205000*60%= 123000
February sales revenue 205000*40%= 82000
March sales revenue 200000*60%=120000
March sales revenue 200000*40%=80000
April sales revenue 190000*60%=114000
Total cash available 283000 231000 237500
Less- Cash payments
Payment of merchandise 300000*70%=210000 205000*70%=143500 200000*70%=140000
Operating expenses 44000 44000 44000
Closing cash balance 29000 43500 53500